OpenStudy (anonymous):

Write an expression for the volume of a cylinder with a height of 7in. greater than the radius

OpenStudy (anonymous):

Well volume of a cylinder is \[\large V=\pi*r^{2}*h\] and we're saying that h=r+7, so just put "r+7" instead of "h".

OpenStudy (anonymous):

is that all you have too do?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Thank you

OpenStudy (pawanyadav):

V=Π×r^2 ×(r+7)

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]